# A polynomial delay algorithm generating all potential maximal cliques in triconnected planar graphs

### v2

@@ 3 Characterizing PMCs in triconnected planar graphs
We show that $X$ satisfies the three conditions for PMCs: (1) There is no full component associated with $X$; (2) Every member of ${\mathcal{C}}_{G}(X)$ is minimally separated; (3) For every pair of vertices $x$ and $y$ in $X$, either $x$ is adjacent to $y$ in $G$ or there is some $C\in{\mathcal{C}}_{G}(X)$ such that $x,y\in N_{G}(C)$.
Condition (1) holds, since each member of ${\mathcal{C}}_{G}(X)$ is contained in one of the regions bounded by cycles $\gamma\in\Gamma\cup\{\gamma_{0}\}$ (Proposition 2.9) and $V(\gamma)$ is a proper subset of $X$ for each such $\gamma$. For Condition (2), let $C$ be an arbitrary member of ${\mathcal{C}}_{G}(X)$. Then, the region of $L_{G}[X]$ that contains $C$ is bounded by a chordless cycle $\gamma\in\Gamma\cup\{\gamma_{0}\}$ of $L_{G}$. Since one side of this cycle contains $C$ and the other side is also non-empty because it contains some vertices of $X$, $V(\gamma)$ is a minimal separator of $G$ due to Proposition 2.11. Therefore, $C$ is minimally separated.
To show that Condition (3) holds, we first show that, for arbitrary two members $x$ and $y$ of $X$, there is some $\gamma\in\Gamma\cup\{\gamma_{0}\}$ such that $x,y\in V(\gamma)$. If $x,y\in V(\gamma_{0})$ then this is trivial so let $x\in X\setminus V(\gamma_{0})$ and let $y\in X$ be arbitrary. First suppose that $P$ is a singleton $\{u\}$, $x=u$, and $y\in V(\gamma_{0})$. Then, we have $x,y\in V(\gamma)$ for some $\gamma\in\Gamma$, since every $v\in S$ belongs to $V(\gamma)$ for some $\gamma\in\Gamma$ and $u$ belongs to $V(\gamma)$ for every $\gamma\in\Gamma$. Next suppose that $|P|\geq 2$. Let $u_{1}$ and $u_{2}$ be two ends of $L_{G}[P]$ and let $R_{i}=N_{L_{G}}(u_{i})\cap S$ for $i\in\{1,2\}$. Due to the definition of a steering, $\{R_{1},R_{2}\}$ is a socket of $\gamma_{0}$. Let $I=P\setminus\{u_{1},u_{2}\}$ be the set of internal vertices of the path $L_{G}[P]$. Due to the definition of a steering, $N_{L_{G}}(I)\cap S$ is either empty or a singleton $\{s\}$ where $s$ is a hinge of the socket $\{R_{1},R_{2}\}$. Suppose first that $N_{L_{G}}(I)\cap S=\emptyset$. Then, the two slots $R_{1}$ and $R_{2}$ divide the cycle $\gamma_{0}$ into two subpaths $p_{1}$ and $p_{2}$ each between a vertex in $R_{1}$ and a vertex in $R_{2}$ with all internal vertices belonging to $V(\gamma_{0})\setminus(R_{1}\cup R_{2})$. Since $R_{1}$ and $R_{2}$ are slots of $\gamma_{0}$, we have $V(p_{1})\cup V(p_{2})=V(\gamma_{0})$. We have a cycle $\gamma_{i}\in\Gamma$ formed by $p_{i}$ and $L_{G}[P]$ together with two edges from $E$, for $i\in\{1,2\}$. Since $S\subseteq V(\gamma_{1})\cup V(\gamma_{2})$ and $P\subseteq V(\gamma_{1})\cap V(\gamma_{2})$, for every pair $(x,y)$ of members of $X$, either $x,y\in V(\gamma_{0})$, $x,y\in V(\gamma_{1})$, or $x,y\in V(\gamma_{2})$.
We turn to the case where $N_{L_{G}}(I)\cap S=\{s\}$, where $s$ is a hinge of the socket $\{R_{1},R_{2}\}$. Let $p$ be the subpath of the cycle $\gamma_{0}$ induced by $V(\gamma_{0})\setminus\{s\}$. Since $R_{1}\cup R_{2}$ is not a slot, one end of $p$ is in $R_{1}$, the other end is in $R_{2}$, and no internal vertex of $p$ is in $R_{1}$ or $R_{2}$. This path $p$ and the path $L_{G}[P]$, together with two edges in $E$, form a cycle $\gamma_{1}\in\Gamma$. If we do not have both $x$ and $y$ in $\gamma_{1}$, then one of $x$ and $y$, say $x$, equals $s$. Since we are assuming that not both of $x$ and $y$ are in $V(\gamma_{0})$, we have $y\in P$. Since every $\gamma\in\Gamma\setminus\{\gamma_{1}\}$ contains $s$ and every vertex in $P$ is contained in some $\gamma\in\Gamma\setminus\{\gamma_{1}\}$, it follows that $x,y\in\gamma$ for some $\gamma\in\Gamma\setminus\{\gamma_{1}\}$. We conclude that, for every pair of two members $x$ and $y$ of $X$, there is some $\gamma\in\Gamma\cup\{\gamma_{0}\}$ such that $x,y\in V(\gamma)$.
@@
We are ready to show that Condition (3) holds. Let $x$ and $y$ be arbitrary two vertices in $X$. As we have shown above, there is a cycle $\gamma$ bounding a region of $L_{G}[X]$ such that $x,y\in V(\gamma)$. If this region contains some $C\in{\mathcal{C}}_{G}(X)$, then we are done since $N_{G}(C)=V(\gamma)$. Otherwise, $\gamma$ is a triangle bounding an empty region and hence $\{x,y\}$ is an edge of $L_{G}[X]$. If this is an edge of $G$ then we are done, so suppose not. Let $\gamma^{\prime}$ be another cycle bounding a region of $L_{G}[X]$ that contains the edge $\{x,y\}$. This region bounded by $\gamma^{\prime}$ cannot be empty since, if it were, then we would have a face of $G$ incident to four or more vertices of $X$, a contradiction to the assumption that $L_{G}[X]$ is a steering and hence is a plane graph. We are done, since $x,y\in N_{G}(C^{\prime})$ where $C^{\prime}$ is the member of ${\mathcal{C}}_{G}(X)$ contained in the region bounded by $\gamma^{\prime}$.
@@
Proof 3.10.
Recall that, due to Proposition 2.16, $L_{G}[X]$ is a biconnected plane graph. Recall also that every $C\in{\mathcal{C}}_{G}(X)$ is contained in a region of $L_{G}[X]$ bounded by a cycle whose vertex set is $N_{G}(C)$. Let $\gamma$ be a cycle that bounds a region $r$ of $L_{G}[X]$ and let $S=V(\gamma)$. If $r$ is empty then $|S|=3$ since otherwise $L_{G}[X]$ would not be planar (Proposition 2.7). Otherwise $r$ contains some $C\in{\mathcal{C}}_{G}(X)$ and we also have $|S|=3$ due to the assumption. Let $x\in X\setminus S$ and $s\in S$ be arbitrary. We show that $x$ is adjacent to $s$ in $L_{G}[X]$. If $x$ is adjacent to $s$ in $G$ then we are done. Otherwise, due to the third condition for PMCs, there is some $C\in{\mathcal{C}}_{G}(X)$ such that $x,s\in N_{G}(C)$. Since the region of $L_{G}[X]$ containing $C$ is bounded by a triangle on $N_{G}(C)$, $x$ is adjacent to $s$ in $L_{G}[X]$. Therefore, $x$ is adjacent in $L_{G}[X]$ to every $s\in S$ and hence $L_{G}[S\cup\{x\}]$ is a $K_{4}$. As $L_{G}[S\cup\{x\}]$ is an $(S,\{x\})$-steering, we have $X=S\cup\{x\}$ due to Proposition 3.7 and hence $L_{G}[X]$ is a $K_{4}$.
@@
Let $G$ be a triconnected plane graph and let $S$ be a minimal separator of $G$. An arch of $S$ is a subset $P$ of $V(G)\setminus S$ such that $L_{G}[P]$ is connected and $N_{L_{G}}(P)\cap S$ is neither empty nor a slot of the cycle $L_{G}[S]$.
@@
Let $S$ be a minimal separator of a triconnected plane graph $G$ with $|S|\geq 4$. Suppose $S$ has an arch and let $P$ be an inclusion-wise minimal arch of $S$. Then, $L_{G}[P\cup S]$ is an $(S,P)$-steering.
Proof 3.13.
Let $H=L_{G}[P\cup S]$. We verify the conditions for $H$ being an $(S,P)$-steering in Definition 3.1. The first condition that $N_{H}(P)$ is neither empty nor a slot of $H[S]$ is satisfied because $N_{H}(P)=N_{L_{G}}(P)\cap S$, $H[S]=L_{G}[S]$, and $P$ is an arch of $S$. If $|P|\geq 2$, $H$ must satisfy the following additional conditions.
  1. 1.
    $H[P]$ is a path. Let $u_{1}$ and $u_{2}$ be the two ends of $H[P]$.
  2. 2.
    The unordered pair $\{R_{1},R_{2}\}$, where $R_{1}=N_{H}(u_{1})\cap S$ and $R_{2}=N_{H}(u_{2})\cap S$, is a socket of $H[S]$.
  3. 3.
    $N_{H}(P\setminus\{u_{1},u_{2}\})\cap S$ is either empty or a singleton $\{s\}$ where $s$ is a hinge of the socket $\{R_{1},R_{2}\}$.
Since $P$ is an arch of $L_{G}[S]$, there are distinct $w_{1},w_{2}\in N_{L_{G}}(P)\cap S$ that are not adjacent to each other in $L_{G}[S]$. Let $u_{i}\in P$ be adjacent in $L_{G}$ to $w_{i}$, for $i\in\{1,2\}$, with $u_{1}$ and $u_{2}$ not necessarily distinct. Let $p$ be a path of $L_{G}[P]$ between $u_{1}$ and $u_{2}$. Then, $V(p)$ is an arch of $L_{G}[S]$ and, since $P$ is an inclusion-wise minimal arch, $L_{G}[P]$ is the path $p$ with ends $u_{1}$ and $u_{2}$. If $u_{1}=u_{2}$ then $P$ is a singleton and $H$ is an $(S,P)$-steering. So, suppose $u_{1}\neq u_{2}$ and hence $|P|\geq 2$. Then Condition 1 holds. For Condition 2, first observe that $R_{i}=N_{H}(u_{i})\cap S=N_{L_{G}}(u_{i})\cap S$ is a slot of $H[S]=L_{G}[S]$, for $i\in\{1,2\}$. This is because, if $R_{1}$ is not a slot of $L_{G}[S]$, say, then $\{u_{1}\}$ would be an arch of $S$, contradicting the minimality of $P$. Since $P$ is an arch of $S$, $N_{L_{G}}(P)\cap S$ is not a slot of $L_{G}[S]$ and hence contains two vertices $y_{1}$ and $y_{2}$ not adjacent to each other on $L_{G}[S]$. Unless $y_{1}$ belongs to one of $R_{1}$ and $R_{2}$ and $y_{2}$ belongs to the other, either $N_{L_{G}}(P\setminus\{u_{1}\})$ or $N_{G_{L}}(P\setminus\{u_{2}\})$ contains both $y_{1}$ and $y_{2}$ and hence there is a proper subset of $P$ that is an arch of $S$, contradicting the minimality of $P$. Therefore, $R_{1}\cup R_{2}$ is not a slot of $L_{G}[S]$ and hence $\{R_{1},R_{2}\}$ is a socket of $L_{G}[S]$.
To show that Condition 3 holds, suppose that $N_{H}(P\setminus\{u_{1},u_{2}\})\cap S$ has a member $s$ that is not a hinge of the socket $\{R_{1},R_{2}\}$. Then, not both of $R_{1}\cup\{s\}$ and $R_{2}\cup\{s\}$ are slots of $L_{G}[S]$. Therefore, either $P\setminus\{u_{1}\}$ or $P\setminus\{u_{2}\}$ would be an arch of $S$, contradicting the minimality of $P$. It follows that every member of $N_{L_{G}}(P\setminus\{u_{1},u_{2}\})\cap S$ is a hinge of the socket $\{R_{1},R_{2}\}$. Even though there can be two hinges $s_{1}$ and $s_{2}$ of the socket when $|S|=4$, $N_{L_{G}}(P\setminus\{u_{1},u_{2}\})\cap S$ cannot contain both of them, since if they did then $P\setminus\{u_{1},u_{2}\}$ would be an arch of $S$, a contradiction. Therefore, Condition 3 holds.
@@
Let $X$ be a PMC of a triconnected plane graph $G$ and let $S\subset X$ be a minimal separator of $G$. Then, there is an arch $P$ of $S$ such that $P\subseteq X\setminus S$.
Proof 3.15.
Since a PMC does not have a full component associated with it, $X\setminus S$ is non-empty. Note also that $X\setminus S$ is contained in one of the two regions of $L_{G}[S]$, since any minimal separator of $G$ that is a subset of a PMC of $G$ does not separate the PMC. Call this region $r_{0}$.
For each edge $e$ of the cycle $L_{G}[S]$, we define a path $p_{e}$ of $L_{G}[X]$ as follows. Consider paths of $L_{G}[X]$ between the ends of $e$ with all internal vertices belonging to $X\setminus S$ and choose such a path $p_{e}$ that maximizes the area in $r_{0}$ bounded by the cycle formed by $p_{e}$ with $e$. As a special case, if the only such path is a path consisting only of $e$, we let $p_{e}$ be this single-edge path. Since $X\setminus S$ is not empty and $L_{G}[X]$ is biconnected (Proposition 2.16), there is some edge $e$ of $L_{G}[S]$ such that $p_{e}$ is not a single-edge path. Let $\gamma$ be the closed walk obtained by concatenating the paths $p_{e}$ for all edges $e$ of the cycle $L_{G}[S]$.
We observe that $\gamma$ cannot be a cycle bounding a region of $L_{G}[X]$, since if it was then it must be chordless due to Proposition 2.16, while an edge $e$ of $L_{G}[S]$ such that $p_{e}$ is not a single edge path is a chord of $\gamma$, a contradiction.
Since $\gamma$ is not a cycle bounding a region of $L_{G}[X]$, either a vertex $v$ repeats on $\gamma$ or $\gamma$ is a cycle but has a chord in the face $r_{1}$ of $\gamma$ that is entirely contained in $r_{0}$. If a vertex $v$ repeats on $\gamma$ then there are two edges $e,e^{\prime}$ of $L_{G}[S]$ such that $v$ is an internal vertex of both $p_{e}$ and $p_{e^{\prime}}$. We claim that if $\gamma$ is a cycle and has a chord $\{x,y\}$ in $L_{G}$ that lies in $r_{1}$ then either (1) $x$ and $y$ are internal vertices of $p_{e}$ and $p_{e^{\prime}}$, respectively, for some distinct edges $e,e^{\prime}$ of $L_{G}[S]$ or (2) one of $x$ and $y$, say $x$, belongs to $S$ and the other, $y$, is an internal vertex of $p_{e}$ for some edge $e$ of $L_{G}[X]$ such that $x\not\in V(e)$. This is because an edge $\{x,y\}$ with $x,y\in S$ is an edge of $L_{G}[S]$ and cannot lie in $r_{1}$ and an edge $\{x,y\}$ with $x,y\in V(p_{e})$ for some edge $e$ of $L_{G}[S]$ cannot lie in $r_{1}$ because of the choice of $p_{e}$ to maximize the area in $r_{0}$ bounded by $p_{e}$ and $e$.
In case (1), as well as when a vertex $v$ repeats on $\gamma$, we let $P=(V(p_{e})\cup V(p_{e^{\prime}}))\setminus((V(e)\cup V(e^{\prime}))$, where $e$ and $e^{\prime}$ are edges of $L_{G}[S]$ specified above in the respective cases. In case (2), we let $P=V(p_{e})\setminus V(e)$. In either case, $P$ is an arch of $S$ since it is connected and we have $|N_{L_{G}}(P)\cap S|\geq 3$.

### v3

@@ 3 Characterizing PMCs in triconnected planar graphs
We show that $X$ satisfies the two conditions for PMCs in Lemma 2.2: (1) There is no full component associated with $X$; (2) For every pair of vertices $x$ and $y$ in $X$, either $x$ is adjacent to $y$ in $G$ or there is some $C\in{\mathcal{C}}_{G}(X)$ such that $x,y\in N_{G}(C)$.
To show that Condition (1) holds, let $C$ be an arbitrary component in ${\mathcal{C}}_{G}(X)$ and $\gamma_{C}$ be the cycle bounding the region of $H$ containing $C$. If $\gamma_{C}$ is $H[S]$, then $C$ is not a full component associated with $X$ as $S$ is a proper subset of $X$. Otherwise, there is a vertex in $S$ that does not belong to $\gamma_{C}$, since $N_{H}(P)$ is not a slot of the cycle $H[S]$, and therefore $C$ is not a full component associated with $X$.
To show that Condition (2) holds, we first show that, for arbitrary two members $x$ and $y$ of $X$, there is some region of $H$ whose boundary contains both $x$ and $y$. If $x,y\in S$, then this certainly holds since $H$ has a region bounded by the cycle $H[S]$. So suppose that at least one of $x$ and $y$, say $x$, belongs to $P$. We argue in a few cases.
First suppose that $|P|=1$ and hence $P=\{x\}$ and $y\in S$. Then, there is certainly a region of $H$ bounded by a cycle consisting of $x$ and a subpath of the cycle $H[S]$ that contains $y$. See the first three examples in Figure 2.
Next suppose that $|P|\geq 2$. Since there is no internal vertex of the path $H[P]$ adjacent in $H$ to any vertex in $S$, there are two regions of $H$ bounded by cycles containing the path $H[P]$. Let $\gamma_{1}$ and $\gamma_{2}$ be those cycles. Then, $\gamma_{i}$, for each $i\in\{1,2\}$, consists of $H[P]$ and a subpath $p_{i}$ of $H[S]$. Since $N_{H}(t)\cap S$ for each end $t$ of the path $H[P]$ is a slot, we have $V(p_{1})\cap V(p_{2})=S$. Therefore, either $\gamma_{1}$ or $\gamma_{2}$ contains both $x$ and $y$. See the last three examples in Figure 2.
@@
We are ready to show that Condition (2) holds. Let $x$ and $y$ be arbitrary two vertices in $X$. As we have shown above, there is a cycle $\gamma$ bounding a region of $H$ such that $x,y\in V(\gamma)$. If this region contains some $C\in{\mathcal{C}}_{G}(X)$, then we are done since $N_{G}(C)=V(\gamma)$. Otherwise, $\gamma$ bounds an empty region and must be a triangle since otherwise $L_{G}[X]$ would not be a plane graph. Therefore, $\{x,y\}$ is an edge of $L_{G}[X]$. If this is an edge of $G$ then we are done, so suppose not. Let $\gamma^{\prime}$ be another cycle bounding a region of $L_{G}[X]$ that contains the edge $\{x,y\}$. This region bounded by $\gamma^{\prime}$ cannot be empty since, if it were, then we would have a face of $G$ incident to four or more vertices of $X$, and $L_{G}[X]$ would not be a plane graph. We are done, since $x,y\in N_{G}(C^{\prime})$ where $C^{\prime}$ is the member of ${\mathcal{C}}_{G}(X)$ contained in the region bounded by $\gamma^{\prime}$.
The converse of this lemma is shown in several steps. We start with a technical proposition.
@@
The next lemma deals with a special case where every minimal separator contained in a PMC $X$ consists of three vertices.
@@
Proof 3.9.
Recall that, due to Proposition 2.17, $L_{G}[X]$ is a biconnected plane graph. Recall also that every $C\in{\mathcal{C}}_{G}(X)$ is contained in a region of $L_{G}[X]$ bounded by a cycle whose vertex set is $N_{G}(C)$. Let $\gamma$ be a cycle that bounds a region $r$ of $L_{G}[X]$ and let $S=V(\gamma)$. If $r$ is empty then $|S|=3$ since otherwise $L_{G}[X]$ would not be a plane graph (Proposition 2.8). Otherwise $r$ contains some $C\in{\mathcal{C}}_{G}(X)$ and we also have $|S|=3$ due to the assumption. Let $x\in X\setminus S$ and $s\in S$ be arbitrary. We show that $x$ is adjacent to $s$ in $L_{G}[X]$. If $x$ is adjacent to $s$ in $G$ then we are done. Otherwise, due to the second condition for PMCs in Lemma 2.2, there is some $C\in{\mathcal{C}}_{G}(X)$ such that $x,s\in N_{G}(C)$. Since the region of $L_{G}[X]$ containing $C$ is bounded by a triangle on $N_{G}(C)$, $x$ is adjacent to $s$ in $L_{G}[X]$. Therefore, $x$ is adjacent in $L_{G}[X]$ to every $s\in S$ and hence $L_{G}[S\cup\{x\}]$ is a $K_{4}$. As $L_{G}[S\cup\{x\}]$ is an $(S,\{x\})$-steering, we have $X=S\cup\{x\}$ due to Proposition 3.6 and hence $L_{G}[X]$ is a $K_{4}$.
To deal with the more general case, we use the following notion of arches.
@@
Let $G$ be a triconnected plane graph and let $S$ be a minimal separator of $G$ with $|S|\geq 4$. An arch of $S$ is a subset $P$ of $V(G)\setminus S$ such that $L_{G}[P]$ is a path and $N_{L_{G}}(P)\cap S$ is neither empty nor a slot of the cycle $L_{G}[S]$.
If $L_{G}[S\cup P]$ is an $(S,P)$-steering, then $P$ is certainly an arch of $S$. The converse does not hold: if $P$ is an arch of $S$ then $L_{G}[S\cup P]$ is not necessarily an $(S,P)$-steering, as the definition of steerings requires more conditions to be satisfied by $L_{G}[S\cup P]$. We show, however, in the next lemma that if $P$ is an inclusion-wise minimal arch then $L_{G}[S\cup P]$ is a steering. A caveat: $L_{G}[S\cup P]$ is not necessarily an $(S,P)$-steering in this case; it may be an $(S^{\prime},\{s\})$-steering where $S^{\prime}=(S\cup P)\setminus\{s\}$ for some $s\in S$. In Lemma 3.13, we show that if $X$ is a PMC then every minimal separator $S\subset X$ with $|S|\geq 4$ has an arch $P$ that is a subset of $X\setminus S$. It follows from these two lemmas that, for every minimal separator $S$ with $|S|\geq 4$ that is a subset of a PMC $X$, there is some $P\subset X\setminus S$ such that $L_{G}[S\cup P]$ is a steering. Due to Lemma 3.4, $S\cup P$ is a PMC and, since no proper subset of a PMC is a PMC (Lemma 2.4), we conclude that $X=S\cup P$ and $L_{G}[X]$ is a PMC, which is an essential part of Theorem 3.15.
@@
Let $S$ be a minimal separator of a triconnected plane graph $G$ with $|S|\geq 4$. Suppose $S$ has an arch and let $P$ be an inclusion-wise minimal arch of $S$. Let $H=L_{G}[S\cup P]$. Then, either $H$ is an $(S,P)$-steering or there is some $s\in N_{L_{G}}(P)\cap S$ such that $H$ is an $((S\cup P)\setminus\{s\},\{s\})$-steering.
Proof 3.12.
If there is some $v\in P$ such that $N_{L_{G}}(v)\cap S$ is neither empty nor a slot of the cycle $L_{G}[S]$, then $\{v\}$ is an arch of $S$ and, since $P$ is minimal, $P=\{v\}$. We are done, since $H$ is an $(S,\{v\})$-steering.
Suppose $N_{L_{G}}(v)\cap S$ is either empty or a slot of $L_{G}[S]$ for every $v\in P$. Since $N_{L_{G}}(P)\cap S$ is not a slot of $L_{G}[S]$, there are two vertices $v_{1},v_{2}\in P$ and two vertices $s_{1},s_{2}\in S$ such that $v_{i}$ is adjacent to $s_{i}$ for $i\in\{1,2\}$ and $\{s_{1},s_{2}\}$ is not an edge of $S$. Let $p$ be a shortest path in $L_{G}[P]$ between $v_{1}$ and $v_{2}$. Then, $V(p)$ is an arch of $S$ and, since $P$ is minimal, $P$ must be equal to $V(p)$. If no internal vertex of $p$ is adjacent in $L_{G}$ to any vertex in $S$, then $L_{G}[S\cup P]$ is an $(S,P)$-steering and we are done. So suppose an internal vertex $v$ of $p$ is adjacent to some vertex in $S$. Let $p_{i}$ be the subpath of $p$ between $v_{i}$ and $v$, for $i\in\{1,2\}$. Since $P$ is a minimal arch of $S$, neither $V(p_{1})$ nor $V(p_{2})$ is an arch of $S$. Therefore, $N_{L_{G}}(\{v_{i},v\})\cap S$ is a slot for $i\in\{1,2\}$. This is possible only if there is a vertex $s\in S$ such that $s$ is adjacent to both $s_{1}$ and $s_{2}$ on $L_{G}[S]$, $N_{L_{G}}(v_{i})$ is either $\{s_{i}\}$ or $\{s_{i},s\}$ for $i\in\{1,2\}$, and $N_{L_{G}}(v)\cap S=\{s\}$. Moreover, since $P=V(p)$ is a minimal arch of $S$, no vertices in $P$ other than $v$, $v_{1}$, or $v_{2}$ are adjacent to any vertex in $S$. Let $p_{3}$ be the subpath of $L_{G}[S]$ between $s_{1}$ and $s_{2}$ avoiding $s$ and let $\gamma$ be the cycle consisting of $p_{3}$ and $p$. Let $S^{\prime}=V(\gamma)=(S\cup P)\setminus\{s\}$. We claim that $\gamma$ is chordless in $L_{G}$, that is, $L_{G}[S^{\prime}]=\gamma$. This is because $p_{3}$ is chordless since $L_{G}[S]$ is a cycle by assumption, $p$ is chordless since it is the shortest path in $P$ between $v_{1}$ and $v_{2}$, and the only edges between $V(p)$ and $V(p_{3})$ are the edges $\{v_{1},s_{1}\}$ and $\{v_{2},s_{2}\}$. Since $N_{L_{G}}(s)\cap S^{\prime}=\{s_{1},v,s_{2}\}$, $L_{G}[S\cup P]$ is an $(S^{\prime},\{s\})$-steering as desired.
@@
Let $X$ be a PMC of a triconnected plane graph $G$ and let $S\subseteq X$ be a minimal separator of $G$ such that $|S|\geq 4$. Then, there is an arch $P$ of $S$ such that $P\subseteq X\setminus S$.
Proof 3.14.
Due to Lemma˜2.1, $X\setminus S$ is non-empty and is a subset of one of the two full components associated with $S$. Let $r_{0}$ be the region of $L_{G}[S]$ that contains $X\setminus S$.
Suppose that there is no arch $P$ of $S$ such that $P\subseteq X\setminus S$. Then, for every pair of vertices $s$ and $s^{\prime}$ in $S$ that are not adjacent to each other on the cycle $L_{G}[S]$, we can draw a curve between $s$ and $s^{\prime}$ in $r_{0}$ without crossing the drawing of $L_{G}[X]$. Therefore, there is a region $r$ of $L_{G}[X]$ contained in $r_{0}$ that is incident to all vertices in $S$. Let $\gamma$ be the cycle of $L_{G}[X]$ bounding $r$. Since $|V(\gamma)|\geq|S|\geq 4$, $r$ is non-empty (Proposition 2.10) and hence contains a component $C\in{\mathcal{C}}_{G}(X)$. Therefore, $V(\gamma)$ must be a minimal separator of $G$, due to Lemma 2.1. But this is impossible, since some edge of $L_{G}[S]$ is a chord of $\gamma$, contradicting Proposition 2.12.